OpenStudy (anonymous):

please help me out in this ... integrate this equation sin^4(e^2x)cos(e^2x)e^2xdx. thanks

OpenStudy (anonymous):

use substitution u=e^2x du=2e^2xdx 1/2du=e^2xdx U will get 1/2 sin^4(u)cos^4(u) du 1/2 (sinucosu)^4 1//2(1/2sin2u)^4 1/32(sin2u)^4

OpenStudy (anonymous):

is it clear until this?

OpenStudy (anonymous):

thank you so much...

OpenStudy (anonymous):

can i ask again?? bout the integration of e^4u du/(1+3e^4u)^2

OpenStudy (anonymous):

use substitution w=1+3e^4u than dw=12e^4u du 1/12dw=e^4udu

OpenStudy (anonymous):

so it will be 1/12 1/w^2 dw integrate to get -1/12w

OpenStudy (anonymous):

and plug w=1+3e^4u back into this

OpenStudy (anonymous):

i get it thanks.